### v1

# DAASH: A Meta-Attack Framework for Synthesizing Effective and Stealthy Adversarial Examples

Abstract
Numerous techniques have been proposed for generating adversarial examples in white-box settings under strict $\ell_{p}$-norm constraints. However, such norm-bounded examples often fail to align well with human perception, and only recently have a few methods begun specifically exploring perceptually aligned adversarial examples. Moreover, it remains unclear whether insights from $\ell_{p}$-constrained attacks can be effectively leveraged to improve perceptual efficacy. In this paper, we introduce DAASH, a fully differentiable meta-attack framework that generates effective and perceptually aligned adversarial examples by strategically composing existing $\ell_{p}$-based attack methods. DAASH operates in a multi-stage fashion: at each stage, it aggregates candidate adversarial examples from multiple base attacks using learned, adaptive weights and propagates the result to the next stage. A novel meta-loss function guides this process by jointly minimizing misclassification loss and perceptual distortion, enabling the framework to dynamically modulate the contribution of each base attack throughout the stages. We evaluate DAASH on adversarially trained models across CIFAR-10, CIFAR-100, and ImageNet. Despite relying solely on $\ell_{p}$-constrained based methods, DAASH significantly outperforms state-of-the-art perceptual attacks such as AdvAD—achieving higher attack success rates (e.g., 20.63% improvement) and superior visual quality, as measured by SSIM, LPIPS, and FID (improvements of $\approx 11$, 0.015, and 5.7, respectively). Furthermore, DAASH generalizes well to unseen defenses, making it a practical and strong baseline for evaluating robustness without requiring handcrafted adaptive attacks for each new defense.
1 Introduction
Adversarial examples against deep neural networks (DNNs), which add imperceptible perturbations to clean images to induce misclassification (targeted/untargeted) Goodfellow et al. (2015a), have attracted significant attention from the research community. Various attack methods, including FGSM Goodfellow et al. (2015b), PGD Madry et al. (2018b), and CW Carlini and Wagner (2017), as well as adaptive evaluation frameworks such as AutoAttack Croce et al. (2020), have been developed to benchmark both defended and undefended models. However, the majority of these approaches prioritize imperceptibility by constraining perturbations within a small $\ell_{p}$-norm ball, with $\ell_{2}$ and $\ell_{\infty}$ norms being the most common ones.
Yet, minimizing $\ell_{p}$-norm does not always yield high perceptual similarity to the human visual system Zhang et al. (2018); Laidlaw and Feizi (2020); Sharif et al. (2018), despite imperceptibility being a fundamental design goal of stealthy adversarial examples. This disconnect has motivated the adoption of perceptually-aligned similarity metrics such as Structural Similarity Index (SSIM) Wang et al. (2004), Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018), and Fréchet Inception Distance (FID) Heusel et al. (2017) to better assess visual quality Li et al. (2024); Chen et al. (2024b). However, attacks optimized under these perceptual metrics often suffer from reduced success rates—particularly against robust, defended models (see Table 1)—highlighting a critical trade-off and leaving significant room for improvement.
Designing effective attacks against a wide range of models, especially hardened ones, remains non-trivial and often requires deep, manual effort—an observation consistently echoed in the literature on $\ell_{p}$-norm attacks Carlini and Wagner (2017); Athalye et al. (2018); Croce and Hein (2020b). This complexity intensifies when attempting to simultaneously preserve high perceptual fidelity, thus underscoring the need for systematic and automated strategies that can generalize across models while aligning better with human vision.
Figure 1: Comparative visualization of adversarial perturbations across multiple attack methods (including DAASH).
Given the extensive body of work on $\ell_{p}$-bounded attacks, including those targeting robust models, it is natural to ask:
Can we leverage the strengths of these existing attacks to compose stronger adversarial examples that are also more perceptually aligned?
Our key insight is that although individual $\ell_{p}$-norms (e.g., $\ell_{2}$, $\ell_{\infty}$) do not fully capture human perception, they each correlate with different perceptual characteristics (e.g., contrast, texture, edge disruption) Sharif et al. (2018); Laidlaw and Feizi (2020). Therefore, by combining attack strategies optimized under different norms in a principled manner, we may achieve not only improved attack success but also better perceptual quality, particularly when targeting defended models where individual attack methods may fall short. Specifically, we propose DAASH (Differential Adversarial Attack SearcH)—a fully differentiable, multi-stage framework that learns to compose multiple $\ell_{p}$-norm–constrained base attacks using soft attention. DAASH operates in two stages: (1) it computes a soft attention distribution over adversarial examples generated from individual candidate attacks (e.g., PGD, CW) to weight and compose the final adversarial example; (2) it iteratively refines the perturbations across multiple stages by feeding the output from one stage as the input to the next. This iterative chaining is designed to escape poor local minima inherent to the non-convex landscape of adversarial optimization.
The attention weights are optimized end-to-end through a meta-loss that jointly considers (i) attack success (via minimizing the confidence score of the ground-truth class) and (ii) perceptual similarity (via maximizing SSIM). This allows the framework to dynamically adjust the influence of each candidate attack based on the target model. Unlike existing methods that are rigid in balance between imperceptibility (from human perception) and attack success, DAASH learns to trade off these objectives adaptively across different models and datasets.
Empirically, DAASH achieves a significant improvement in both attack success rate (ASR) and perceptual fidelity compared to the current state-of-the-art. On robust models trained on CIFAR-100 such as Cui et al. (2024), DAASH reaches an average ASR of 99.97% with an SSIM score of 94.43, significantly outperforming state-of-the-art perceptual attacks such as AdvAD Li et al. (2024) (ASR: 79.14%, SSIM: 83.18) and DiffAttack Chen et al. (2024b) (ASR: 77.24%, SSIM: 91.16). Similar improvements are observed in other experimental settings.
Fig. 1 presents qualitative comparisons, further illustrating the superiority of DAASH in generating perceptually plausible yet highly effective adversarial examples.
Our contributions are summarized as follows:
  • We propose DAASH, a fully differentiable, multi-stage framework that learns to combine $\ell_{p}$-norm–constrained base attacks using soft attention.
  • We introduce a novel meta-loss that decouples attack effectiveness and human perceptual quality from the constraints of individual norm-based attacks, enabling better identification of attack strengths across models.
  • We validate DAASH across multiple datasets (CIFAR-10, CIFAR-100, ImageNet), seven robust models, and four post-processing defenses demonstrating substantial improvements over prior perceptual attacks.
2 Background and Threat Model
Adversarial Examples. Given a deep neural network based classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$, which maps inputs to class probabilities, and a clean sample $(x,y)$ where $x\in\mathcal{X}$ and $y\in\mathcal{Y}$, an untargeted adversarial example $\boldsymbol{x}_{a}$ is defined as:
$\arg\max_{i}f(\boldsymbol{x}_{a})_{i}\neq y\quad\text{subject to}\quad d(\boldsymbol{x},\boldsymbol{x}_{a})\leq\epsilon$ (1)
Imperceptibility constraints. In the equation above, $d(\cdot,\cdot)$ represents a distance metric and $\epsilon$ defines the perturbation budget. $d(\cdot,\cdot)$ measures the imperceptibility of the perturbation, and the commonly used metrics include $\ell_{2}$ norm Carlini and Wagner (2017) as well as the $\ell_{\infty}$-norm Carlini and Wagner (2017). These $\ell_{p}$-norm constraints may not always align with human perception, and some alternative metrics such as the Structural Similarity Index (SSIM) Wang et al. (2004) is used. SSIM is shown to better capture the luminance, contrast, and structural similarities between two images $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ and is computed as below:
$\text{SSIM}(\boldsymbol{x}_{1},\boldsymbol{x}_{2})=\frac{(2\mu_{x_{1}}\mu_{x_{2}}+c_{1})(2\sigma_{x_{1}x_{2}}+c_{2})}{(\mu_{x_{1}}^{2}+\mu_{x_{2}}^{2}+c_{1})(\sigma_{x_{1}}^{2}+\sigma_{x_{2}}^{2}+c_{2})}$ (2)
The terms $\mu_{x_{1}}$ and $\mu_{x_{2}}$ represent the mean pixel intensities of $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, respectively, while $\sigma_{x_{1}}^{2}$ and $\sigma_{x_{2}}^{2}$ denote their variances. The term $\sigma_{x_{1}x_{2}}$ corresponds to the covariance between $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$. The constants $c_{1}$ and $c_{2}$ are used to stabilize the division when the denominator is close to zero.
Besides SSIM, other perceptual similarity metrics that better align with human perception include the Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018) and the Fréchet Inception Distance (FID) Heusel et al. (2017). In this paper, we use SSIM as the primary metric to optimize for imperceptibility, but we find that using LPIPS and FID yields similar results in practice.
Threat model. In this paper, we focus on the white-box setting, where the attacker is assumed to have complete access to $f(\cdot)$, including the model architecture and parameters. In this setting, untargeted adversarial examples, as defined in Eqn. (1), can be generated via backpropagation using well-defined loss functions with respect to the input $\boldsymbol{x}$. The white-box assumption is widely adopted in the extensive literature Carlini and Wagner (2017); Madry et al. (2018a); Croce and Hein (2020b).
3 Related Work
Classical approaches for generating adversarial examples include FGSM Goodfellow et al. (2015a), I-FGSM Kurakin et al. (2016), and PGD Madry et al. (2018a), all of which strictly enforce a perturbation budget $\epsilon$ measured in the $\ell_{\infty}$-norm. In contrast, the CW attack Carlini and Wagner (2017) softly enforces an $\ell_{2}$-norm perturbation constraint by incorporating the perturbation magnitude directly into the optimization objective. Closer to our approach are AutoAttack Croce and Hein (2020b) and Composite Attacks Mao et al. (2021). AutoAttack employs an ensemble of a few selected attack methods, while Composite Attacks combine multiple attacks using genetic algorithms. However, these approaches still operate under conventional $\ell_{p}$-norm constraints and rely on simplistic or heuristic strategies for combining base attacks. In contrast, our proposed method DAASH introduces an optimization-based meta-framework to adaptively compose base attacks for improved attack effectiveness and imperceptibility for human eyes.
Recognizing the limitations of norm-constrained attacks in modeling human perception, recent works have shifted toward leveraging perceptual similarity metrics to guide adversarial example generation. For instance, SSAH Luo et al. (2022) operates in the frequency domain to introduce perturbations in less perceptible high-frequency components, while PerC-AL Zhao et al. (2020) optimizes perceptual color distance, though it relies on manual hyperparameter tuning. More recent approaches exploit diffusion models by injecting perturbations into their embedding spaces, thereby influencing the denoising trajectory to produce adversarial outputs. Representative examples include DiffAttack Chen et al. (2024a), AdvAD Li et al. (2024), and DiffPGD Xue et al. (2023). Despite their innovations, these perceptual attacks are often designed as standalone strategies with limited adaptability, making them less effective across diverse target models. As demonstrated in Table 1, such methods frequently underperform in both attack success and imperceptibility. By contrast, DAASH builds on the extensive literature of norm-constrained attacks—known for their compatibility with a wide range of target models—and combines them in a unified framework that enhances the effectiveness and stealth of the resulting adversarial examples.
4 DAASH Attack Design
Next, we describe the high-level motivation behind the design of DAASH and then introduce the details more formally.
Figure 2: An overview of the proposed Differentiable Adversarial Attack Search (DAASH) methodology.
4.1 Design motivation
Our attack design is motivated by two key observations: (1) Norm-constrained attacks sometimes exhibit complementary strengths and weaknesses Croce and Hein (2020b). So, combining them can yield better performance, especially in under-explored settings of generating adversarial examples beyond $\ell_{p}$-norm constraints; (2) Adversarial example generation is a non-convex optimization process and is prone to getting stuck in poor local optima Madry et al. (2018a).
Regarding Observation 1, the most natural way to combine different attacks is to take a weighted average of the adversarial examples generated by each individual attack. While a linear combination is simple and practical, it introduces two key challenges: (a) How to determine optimal weights for each base attack, and (b) How to ensure that the resulting adversarial example better aligns with human perception rather than just staying within the $\ell_{p}$ constraints.
To address these, we make the combination weights continuous and learnable. We introduce a meta-loss that jointly optimizes for attack effectiveness and perceptual similarity to human vision. This allows the weights to be adjusted through gradient-based optimization, effectively converting norm-constrained base attacks into perceptually aligned attacks—without needing to modify their core mechanisms.
Regarding Observation 2, we mitigate the risk of poor local optima by introducing a multi-stage chaining process. In each stage, we recombine base attacks using newly learned weights and feed the output adversarial example from the previous stage into the next as input. This design allows the attack to iteratively refine the adversarial perturbation. Prior work on transfer-based attacks has shown the benefits of chaining attacks across stages Suya et al. (2020), but typically only use two stages and assign different attacks to each. In contrast, our method allows arbitrary depth and continuously optimized attack combinations across all stages.
4.2 Design Details
Next, we introduce our attack design more formally. We assume a total of $M$ base attacks and $N$ stages (or attack cells), both of which are hyperparameters that the attacker can freely choose to optimize attack performance. The set of base attacks can also be easily expanded as new attack methods become available in the future.
In a given stage $j$, we denote the learnable weight for each base attack as $\alpha_{j,i}$, where $i$ indexes the base attack. The output of the $j$-th attack cell is then defined as:
$\mathcal{A}_{j}(\boldsymbol{x}^{j-1}_{a},y)=\sum_{i=1}^{M}\frac{\exp(\alpha_{j,i})}{\sum_{k=1}^{M}\exp(\alpha_{j,k})}\cdot\boldsymbol{x}_{a}^{j,i}$ (3)
where $\boldsymbol{x}_{a}^{j,i}$ denotes the adversarial example generated by the $i$-th base attack at stage $j$. The softmax operation ensures that the final combination remains within the original $\ell_{p}$-norm constraints, as we aim to preserve the structure of the base attacks without modifying them directly. This linear formulation provides a simple and extensible mechanism for combining base attacks. Future work may explore more advanced combination strategies beyond weighted averaging, or treat individual hyperparameters of base attacks, which are fixed in this work, as additional learnable variables.
After the adversarial example generation in each stage, we apply multi-stage chaining as follows: $\boldsymbol{x}_{a}=\mathcal{A}_{N}(\boldsymbol{x}_{a}^{N-1},y)\circ\cdots\circ\mathcal{A}_{1}(\boldsymbol{x}^{0},y)$ where the composition operator $\circ$ indicates the output of each stage is passed as input to the next. This enables iterative refinement of the adversarial perturbation over multiple stages. While we adopt a simple linear chaining strategy here, more sophisticated chaining mechanisms (e.g., cyclic or attention) are possible and left as future work.
Finally, we define the meta-loss that guides the optimization of the attack weights: $\mathcal{L}_{\text{total}}=\lambda_{\text{asr}}\cdot f_{y}(\boldsymbol{x}_{a})+\lambda_{\text{ssim}}\cdot(1-\text{SSIM}(\boldsymbol{x},\boldsymbol{x}_{a}))$ The first term encourages attack effectiveness by maximizing the confidence drop in the ground-truth class $y$ for the adversarial example $\boldsymbol{x}_{a}$. The second term penalizes perceptual distortion, measured using the Structural Similarity Index Measure (SSIM), which ranges from 0 to 1, with higher values indicating greater similarity. In practice, we observe that substituting SSIM with other perceptual similarity metrics yields comparable results. Our full attack algorithm is summarized in Algorithm 1 and a visual demonstration is also shown in Figure 2.
Algorithm 1 Search for Optimal Attack Sequence $\mathcal{A}_{\Theta}$
1:Randomly Initialize $\boldsymbol{\alpha}$
[…]
3:for $\text{epoch}=1$ to $T$ do
4:  Sample a mini-batch $\{(\boldsymbol{x},y)\}\subset\mathcal{D}$
5:  Generate adversarial examples: $\boldsymbol{x}_{a}\leftarrow\mathcal{A}(\boldsymbol{x},y)$
6:  Feed to target model: $f(\boldsymbol{x}_{a})$
7:  Compute total loss $\mathcal{L}_{\text{total}}$
8:  Update parameters: $\boldsymbol{\alpha}\leftarrow\boldsymbol{\alpha}-\eta\,\nabla_{\boldsymbol{\alpha}}\mathcal{L}_{\text{total}}(\boldsymbol{x}_{a},y)$
9:end for
10:Return attack sequence $\mathcal{A}_{\Theta}$ that minimizes $\mathcal{L}_{\text{total}}$
Table 1: Performance comparison of adversarial attacks across different datasets and RobustBench models. Base is the ASR without applying any post-processing defenses; Avg. is the average of all the ASR values. SSIM values are scaled to [0,100].
Dataset Target Model Attack ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)
Base JPEG TVM Ensemble NRP Avg.
CIFAR-100 Cui et al. (2024) DI-FGSM 65.20 65.49 65.35 65.55 57.03 63.72 92.08 0.0168 54.93
CW 100.00 87.70 96.68 78.61 57.71 84.14 94.38 0.0153 32.49
PI-FGSM++ 64.41 65.78 64.32 65.64 50.86 62.20 91.12 0.0199 58.94
AutoAttack 82.56 81.33 79.51 80.55 59.88 76.77 92.09 0.0167 41.68
DiffPGD 76.20 76.00 75.80 75.60 75.40 75.80 91.15 0.0165 43.50
DiffAttack 77.50 77.20 77.70 77.40 76.40 77.24 91.16 0.0169 42.30
AdvAD 82.20 82.80 81.10 81.70 67.90 79.14 83.18 0.0288 45.67
[…]
Wang et al. (2023) DI-FGSM 66.46 67.01 66.07 66.72 59.86 65.22 91.96 0.0169 55.24
CW 100.00 85.74 95.31 76.95 59.86 83.57 94.55 0.0161 32.19
PI-FGSM++ 66.17 67.21 65.29 66.76 52.30 63.55 91.14 0.0199 57.88
AutoAttack 83.50 82.91 81.05 81.64 62.60 78.34 91.25 0.0189 42.19
DiffPGD 79.10 76.10 74.20 73.80 73.40 75.32 91.50 0.0161 51.80
DiffAttack 76.90 76.60 76.40 76.80 76.30 76.53 90.90 0.0176 42.20
AdvAD 79.90 79.60 79.90 79.50 67.20 77.42 83.07 0.0293 47.29
[…]
(Addepalli et al.) (2022) DI-FGSM 72.09 72.13 71.84 71.74 63.42 70.24 92.05 0.0156 56.69
CW 100.00 87.21 96.48 77.93 60.74 84.47 95.56 0.0113 26.94
PI-FGSM++ 71.50 71.93 70.86 71.15 56.89 68.47 91.14 0.0192 59.94
AutoAttack 83.77 82.70 81.02 80.68 61.33 77.90 93.03 0.0147 39.59
DiffPGD 82.50 82.20 81.80 81.60 80.90 81.80 91.05 0.0168 42.80
DiffAttack 81.60 81.60 81.40 81.40 81.50 81.35 91.26 0.0163 42.15
AdvAD 84.50 84.00 83.70 84.30 70.50 81.60 83.92 0.0291 45.63
[…]
CIFAR-10 (Bartoldson et al.) (2024) DI-FGSM 41.39 42.46 42.27 42.75 35.70 40.91 90.88 0.0200 59.07
CW 97.27 61.62 96.09 60.35 58.59 74.78 95.25 0.0206 55.25
PI-FGSM++ 45.80 48.46 46.72 48.89 23.24 42.62 88.74 0.0278 66.91
AutoAttack 73.24 71.88 72.75 71.19 39.36 65.68 87.51 0.0264 53.05
DiffPGD 53.20 52.40 52.60 52.40 52.80 52.68 91.60 0.0165 54.20
DiffAttack 73.60 73.30 73.60 73.50 72.90 73.38 82.90 0.0369 51.60
AdvAD 61.30 60.70 60.60 61.20 41.10 56.98 83.50 0.0269 49.37
[…]
Cui et al. (2024) DI-FGSM 51.64 52.71 52.36 53.01 45.23 50.99 90.51 0.0206 62.20
CW 99.80 67.68 95.51 65.33 65.33 78.73 95.39 0.0206 51.38
PI-FGSM++ 55.76 57.09 56.05 56.84 31.07 51.36 88.71 0.0272 68.81
AutoAttack 78.42 77.34 78.22 76.46 44.34 70.96 86.88 0.0273 56.71
DiffPGD 76.20 75.80 76.50 76.10 75.40 76.00 79.80 0.0350 60.50
DiffAttack 82.80 82.50 83.10 82.70 82.20 82.66 81.36 0.0404 58.24
AdvAD 66.80 67.20 66.60 66.20 45.50 62.46 83.11 0.0265 50.95
[…]
Wang et al. (2023) DI-FGSM 51.64 52.13 52.09 51.84 44.57 50.45 90.45 0.0203 63.21
CW 98.90 70.21 94.82 66.11 51.76 76.36 95.25 0.0208 52.86
PI-FGSM++ 55.29 57.11 54.61 56.27 29.61 50.58 88.72 0.0265 70.71
AutoAttack 77.93 77.54 77.73 76.27 43.16 70.53 86.83 0.0276 56.75
DiffPGD 75.80 75.40 76.10 75.60 75.10 75.60 79.50 0.0345 60.80
DiffAttack 82.50 82.40 82.50 82.40 82.10 82.38 80.12 0.0429 57.22
AdvAD 67.40 68.40 66.40 67.20 44.10 62.70 82.64 0.0271 52.78
[…]
ImageNet-1k Salman et al. (2020) DI-FGSM 83.90 83.60 83.00 82.60 75.40 81.70 91.42 0.0849 49.39
CW 98.50 88.40 98.20 76.00 63.70 85.02 97.75 0.0263 6.77
PI-FGSM++ 87.10 86.80 85.60 85.10 74.50 83.82 90.54 0.0960 49.57
AutoAttack 94.90 94.30 91.30 90.30 72.60 91.56 92.90 0.0757 21.78
DiffPGD 76.20 69.40 74.80 68.50 72.30 72.24 87.50 0.0892 45.60
DiffAttack 92.50 86.80 88.90 87.90 89.80 89.18 89.17 0.0763 40.78
AdvAD 89.20 87.90 87.70 87.90 86.80 87.90 71.30 0.1949 50.14
[…]
5 Experiments and Results
Next we discuss our experimental setup and evaluations. All the reported results are obtained using NVIDIA A100 GPU.
5.1 Experimental Setup
5.1.1 Datasets and Models
Our experiments are conducted on three commonly used benchmark datasets of different scales: CIFAR-10 Krizhevsky et al. (2014), CIFAR-100 Krizhevsky (2009), and ImageNet Deng et al. (2009). For each dataset, we randomly sample 1,000 test images that are correctly classified by the victim model. Since standard models (without any defenses) are easily broken by existing attacks, we focus our evaluation on robust models curated by RobustBench Croce et al. (2020), a popular benchmark for adversarial robustness.
5.1.2 Post-Processing Defenses
To further test the limits of DAASH and other SOTA attacks in terms of generating potent adversarial examples, we also experiment with post-processing defenses (on top of using RobustBench models). The post-processing defenses include JPEG compression Guo et al. (2017), Total Variation Minimization (TVM) Wang et al. (2018), Neural Representation Purifier (NRP) Naseer et al. (2020), and an ensemble of four methods: bit-depth reduction Xu (2018), JPEG, TVM, and Non-Local Means Denoising (NLM) Buades et al. (2011).
5.1.3 Evaluation Metrics
We use attack success rate (ASR) to measure attack potency, and for determining human visual imperceptibility, we use: SSIM, LPIPS, and FID.
Figure 3: Comparison of DAASH performance before and after training, illustrating ASR and SSIM across different defense settings. “Trained DAASH" significantly outperforms the randomly weighted version (“Random DAASH").
5.1.4 Base Attacks and DAASH Training Details
We include 10 diverse base attacks in our attack pool: FGSM Goodfellow et al. (2015b), PGD Madry et al. (2018b), CW Carlini and Wagner (2017), FAB Croce and Hein (2020a), TI-FGSM Dong (2019), NI-FGSM Lin et al. (2019), MI-FGSM Dong et al. (2017), DI-FGSM Xie et al. (2018), BIM Kurakin et al. (2016), and PI-FGSM++ Gao et al. (2020). Each attack offers unique advantages and covers a broad range of perturbation strategies. In addition, we include a None operation at each stage, which simply returns the input image, allowing the model to flexibly decide whether to skip processing in a given stage.
For the meta-loss, we set $\lambda_{\text{asr}}=1.3$ and $\lambda_{\text{ssim}}=1.0$, which work well across datasets. We run Algorithm 1 for $T=100$ epochs using the Adam optimizer with a learning rate of 0.001. After training, the learned weights for combining base attacks in each stage are used to generate the final adversarial example.
Figure 4: DAASH training progression over epochs.
5.2 Effectiveness of DAASH against Robust Models
Table 1 reports the attack success rate (ASR) and perceptual similarity scores of DAASH and seven representative state-of-the-art attacks evaluated against seven recent adversarially trained robust models across CIFAR-10, CIFAR-100, and ImageNet. As shown in the table, DAASH consistently outperforms all baseline attacks in terms of ASR under both the Base (no post-processing) and post-processing defense conditions, across all datasets. This includes outperforming recent attacks explicitly designed to go beyond $\ell_{p}$-norm constraints. For example, on CIFAR-100, the best-performing baseline is CW with an average ASR of $84\%$, while the most recent perceptual-aligned attack, AdvAD, achieves only $79\%$. In contrast, DAASH achieves an ASR close to $100\%$. Moreover, DAASH yields better imperceptibility scores compared to these dedicated perceptual attacks. Similar trends are observed on CIFAR-10 and ImageNet: DAASH improves over the best baseline by at least $15\%$ ASR on CIFAR-10 and $5\%$ on ImageNet, on average. Its perceptual scores are also consistently better than other attacks specifically designed for perceptual alignment.
Table 2: Effectiveness of DAASH with varying number of stages for the robust CIFAR-100 model – Cui et al. (2024).
Num. of ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)
Stages Base JPEG TVM Ensemble NRP Avg.
1 72.85 34.08 68.46 37.70 46.29 51.88 98.48 0.0027 16.19
2 100.00 99.90 100.00 98.54 99.02 99.49 94.59 0.0115 43.28
3 100.00 99.90 100.00 99.32 99.61 99.77 94.43 0.0139 41.08
4 100.00 100.00 100.00 99.61 99.71 99.86 93.22 0.0190 47.56
5 100.00 100.00 100.00 100.00 99.90 99.98 92.92 0.0209 48.25
Table 3: Average attack success rate (ASR↑) and structural similarity (SSIM↑) for cross-model transferability.
Trained Tested CIFAR-100 CIFAR-10 ImageNet-1K
(Cui et al.) (2024) (Wang et al.) (2023) (Addepalli et al.) (2022) (Bartoldson et al.) (2024) (Cui et al.) (2024) (Wang et al.) (2023) (Salman et al.) (2020)
Dataset ASR SSIM ASR SSIM ASR SSIM ASR SSIM ASR SSIM ASR SSIM ASR SSIM
[…]
Effectiveness of CW. An interesting finding is that the CW attack, originally developed for $\ell_{2}$-norm constrained settings, also achieves competitive ASR and perceptual similarity when its hyperparameters are carefully tuned (we tested multiple configurations and report the best result). In some cases, CW was also able to outperform DiffPGD, AdvAD, and DiffAttack. This suggests that some classical attacks, when properly tuned, can potentially still be highly effective in terms of both ASR and imperceptibility.
Although DAASH outperforms CW in terms of ASR, its imperceptibility scores are sometimes slightly lower. Intuitively, since CW is one of the base attacks in DAASH, it could be expected to receive higher attention weights during optimization. However, this is not always the case. The reason lies in the mismatch between the objectives: CW uses a hinge-like loss that stops contributing gradients once misclassification is achieved, whereas our meta-loss continues to penalize the confidence of the true class beyond the decision boundary. As a result, after initial misclassification, CW receives lower weights in DAASH, which shifts focus to other attacks that still produce informative gradients. This continued “pushing” of adversarial examples beyond the boundary enhances robustness of DAASH against additional defenses such as post-processing but comes at the cost of slightly reduced perceptual similarity.
5.3 Ablation Studies on DAASH
Our DAASH framework consists of two key components: (1) learnable weights for each base attack within every attack cell, and (2) chained attack cells to better escape poor local optima. In this section, we conduct ablation studies to demonstrate the importance of each component.
Importance of Learnable Weights. To evaluate the benefit of using learnable weights, we compare the performance of DAASH against a variant where the weights for base attacks are initialized randomly and left untrained. Specifically, the random weights are sampled uniformly in the range $[0,1]$ and used directly without optimization. We repeat this random setup 100 times and report the average results.
Figure 3 shows that using trained weights significantly improves both ASR and imperceptibility (measured by SSIM) across all types of robust models. Furthermore, Figure 4 illustrates the ASR and SSIM trends as training progresses over epochs. Both metrics exhibit clear improvement—ASR increases across all settings, and SSIM improves almost monotonically—highlighting the effectiveness of our meta-learning process. These results confirm that the learnable weighting mechanism in DAASH plays a crucial role.
Importance of Multiple Stages. To assess the value of chaining multiple stages, we evaluate DAASH with varying numbers of stages. Table 2 reports the ASR for different stage configurations. The results show a substantial ASR jump when increasing from one to two stages (from 51.88% to 99.49%), suggesting that two stages are sufficient to escape bad local optima and refine the adversarial example generation. Adding more stages yields marginal improvements, indicating diminishing returns beyond two stages.
Effectiveness against Unseen Defenses. We also examine whether the learned weights in DAASH are transferable across models and datasets. Table 3 shows that weights trained on one model or dataset can be applied to unseen models or datasets while still achieving near-optimal results. The variation in average ASR remains within 3%, and SSIM varies by less than 2. These results demonstrate that DAASH can generalize well across different settings/models.
6 Conclusion
In this work, we introduced DAASH, a novel differentiable framework for learning optimal compositions of adversarial attacks in a white-box setting. Unlike traditional handcrafted methods that heuristically combine attacks, DAASH leverages gradient-based optimization to learn soft weightings over multiple attack modules across multiple stages. This process is guided by a meta-loss that jointly captures attack effectiveness and perceptual alignment with human vision.
Our flexible framework enables the transformation of existing norm-constrained attacks into stronger attacks that also produce perceptually aligned adversarial examples. Empirical results demonstrate that DAASH  significantly outperforms state-of-the-art perceptually aligned attacks in both attack success rate and imperceptibility scores. DAASH  generalizes well to unseen defenses, establishing itself as a strong and adaptive baseline for future robustness evaluation.
Overall, our results underscore the potential of differentiable search in the adversarial example literature and open up new directions for developing more adaptive and powerful attack strategies. Future work may explore extending DAASH  to black-box or query-limited scenarios, as well as integrating it with generative models to further enhance imperceptibility and increase sample diversity.